OpenStudy (anonymous):

write each expression as the product of two bionomials 2x(y+4)+3(y+4) i don't understand how to figure this out

OpenStudy (aum):

2x(y+4)+3(y+4) 2x(y+4) has the factor (y+4) and 3(y+4) also has the factor (y+4). Pull out the common factor: 2x(y+4)+3(y+4) = (y + 4)(2x + 3)

OpenStudy (anonymous):

so what is the common factor?

OpenStudy (aum):

(y + 4) is the common factor.

OpenStudy (anonymous):

okay thanks i think i understand now

OpenStudy (aum):

You are welcome.

OpenStudy (aum):

What we did was we noticed that in 2x(y+4)+3(y+4) the (y+4) multiplies 2x and then multiplies 3 and adds them. So we can factor out (y+4) and write it as (y+4)(2x + 3).

OpenStudy (aum):

The answer is a product of two binomials. Binomial means TWO terms. y + 4 has two terms and so it is a binomial. 2x + 3 also has two terms and so it is also a binomial

OpenStudy (anonymous):

so you wouldn't use FOIL?

OpenStudy (anonymous):

i actually have another question if you aren't busy

OpenStudy (aum):

You wouldn't use FOIL here because they are not asking you to expand and then simplify. Here they want you to factor 2x(y+4)+3(y+4). But instead of using the word "factor" they are just saying it in a roundabout way as: write it as a product of two binomials. Go ahead and post your other question.

OpenStudy (anonymous):

okay i get that one now... the other question is x^2-x-6 the same directions apply

OpenStudy (aum):

x^2 - x - 6 = (x + ?) * (x + ?) We need to come up with numbers where the question marks are. The two numbers should be such that they add to the coefficient of the middle term (-1) and they should multiply to the constant term (-6). Can you come up with two such numbers? Take the constant term and write all possible pairs of factors for it. Can you list all factors of -6?

OpenStudy (anonymous):

okay so 1,6 2,3 -1,-6 and -2,-3

OpenStudy (aum):

Slight modification: Factors of -6: 1 and -6 -1 and 6 2 and -3 -2 and 3 All of these pairs multiply to -6. But which pair adds to -1?

OpenStudy (anonymous):

2 and -3 correct?

OpenStudy (aum):

Correct. x^2 - x - 6 = (x + 2) * (x - 3) That is it.

OpenStudy (anonymous):

thank you...are you a math teacher?

OpenStudy (aum):

You are welcome. Not a teacher. Just volunteering here to help out with math.

OpenStudy (anonymous):

do you know a lot of algebra 2/ trig because thats what I'm studying and im frankly not very good at math

OpenStudy (aum):

There are many, many people here who are good at those. Just post the questions one at a time and when it is answered satisfactorily you can close that question and post a new one. Also, if a number of people post replies and you like one of the replies the best you can click on the "Best Response" button next to their reply.

OpenStudy (anonymous):

well i guess it's pretty clear that this is my first time on the website